# NeAR: Coupled Neural Asset–Renderer Stack

### v1

@@ Abstract
Neural asset authoring and neural rendering have emerged as fundamentally disjoint threads: one generates digital assets using neural networks for traditional graphics pipelines, while the other develops neural renderers that map conventional assets to images. However, the potential of jointly designing the asset representation and renderer remains largely unexplored. We argue that coupling them can unlock an end-to-end learnable graphics stack with benefits in fidelity, consistency, and efficiency. In this paper, we explore this possibility with NeAR: a Coupled Neural Asset–Renderer Stack. On the asset side, we build on Trellis-style Structured 3D Latents and introduce a lighting-homogenized neural asset: from a casually lit input, a rectified-flow backbone predicts a Lighting-Homogenized SLAT that encodes geometry and intrinsic material cues in a compact, view-agnostic latent. On the renderer side, we design a lighting-aware neural renderer that uses this neural asset, along with explicit view embeddings and HDR environment maps, to achieve real-time, relightable rendering. We validate NeAR on four tasks: (1) G-buffer–based forward rendering, (2) random-lit single-image reconstruction, (3) unknown-lit single-image relighting, and (4) novel-view relighting. Our coupled stack surpasses state-of-the-art baselines in both quantitative metrics and perceptual quality. We hope this coupled asset-renderer perspective inspires future graphics stacks that view neural assets and renderers as co-designed components instead of independent entities.
Figure 1: Relightable 3D generation. Generated assets from casually lit inputs (col. 2) rendered under target lighting (col. 1). Cols. 3-5: traditional Blender rendering (Trellis, Hunyuan3D, ours). Col. 6: our method with neural rendering. Col. 7: GT.
@@ 1 Introduction
Images are determined by the interaction of light with scene geometry, materials, and lighting. Classical computer graphics separates this process into asset authoring, where artists define scene properties, and rendering, where a physically based renderer simulates light transport. While effective, this separation requires substantial manual effort, computationally expensive simulations, and makes inverse reconstruction from real-world images or video challenging. Recent advances in neural graphics [55, 44, 20, 60, 45, 48, 3, 6, 16, 54, 52, 28] address these limitations from two complementary directions: neural asset authoring uses generative models [55, 44, 20, 60, 45, 48, 5, 3, 6] to synthesize full 3D assets for traditional pipelines, reducing manual effort, while neural renderers map these assets—often converted into intermediate representations such as depth, normals, or shading buffers—directly to images [23, 54, 52], providing a data-driven alternative to analytic rendering and enabling more robust inverse inference.
@@
Despite recent progress in generating 3D assets with PBR materials [55, 44, 20, 60, 45, 48, 5, 3, 6], a fundamental limitation remains: asset generation and neural rendering are typically developed in isolation, with assets created assuming a fixed renderer and renderers trained on static asset distributions. This separation becomes problematic when errors in asset decomposition—such as misidentified albedo or incorrect normal maps—propagate through the rendering pipeline. Because rendering is a nonlinear process, small errors in asset decomposition compound into visible artifacts like baked-in shadows or lighting inconsistencies. Fig. 1 demonstrates this issue: existing methods rendered with traditional physically-based renderers (e.g., Blender) exhibit lighting artifacts and fail to achieve faithful relighting.
@@
To this end, we propose NeAR, a Coupled Neural Asset–Renderer Stack for single-image relightable 3D generation. Our key insight is to co-design the asset representation and rendering process to enable relighting directly through a shared, lighting-homogenized latent space. On the asset side, we introduce a Lighting-Homogenized Structured 3D Latent (LH-SLAT). Unlike standard assets that rely on fragile explicit decomposition, our model lifts the casually lit input into a canonical latent form. As visualized in Fig. 2, this process transforms a shadow-affected representation (Shaded-SLAT) into a clean, homogenized state, effectively suppressing baked-in shadows and unstable highlights while preserving geometric cues. On the renderer side, we design a lighting-aware neural renderer. Conditioned on a lighting tokenizer, this renderer learns to interpret the homogenized latents and synthesize view-dependent appearance under arbitrary HDR environments via differentiable 3D Gaussian splatting. By unifying the representation, NeAR generates assets that naturally support real-time, high-quality relighting and novel-view synthesis with consistent materials across views.
Figure 2: Lighting homogenization. Visualization of renders under random lighting and homogenized illumination. Starting from base color and ambient occlusion (AO), we show shaded results under random and canonical lighting, estimated shadows, our lighting-homogenized structured latent (LH-SLAT), and relit outputs compared to ground truth (GT). LH-SLAT suppresses shadows and unstable specularities while preserving geometry-consistent diffuse cues.
@@
    Coupled neural asset–renderer stack. We introduce NeAR, an end-to-end learnable graphics stack where the neural asset representation and neural renderer are co-designed for single-image relightable 3D asset generation.
@@ 2.1 Image relighting and inverse rendering
Image relighting and inverse rendering lie at the intersection of geometry, material estimation, and light transport and have been studied from both physics-driven and data-driven perspectives [16]. Classical inverse rendering methods (e.g., SIRFS) recover interpretable PBR maps (albedo, roughness, normals) via optimization and hand-crafted priors [1]. These pipelines are interpretable and editable, but the inverse problem is highly ill-posed in real scenes: shadows, inter-reflections and view-dependent highlights easily bias the recovered materials and produce baked-in artifacts under re-rendering.
Recent learning-based approaches generally fall into two categories. The first focuses on physically-structured decomposition [60, 23, 10]. Although decomposition yields interpretable assets, accurate regression from casual single-view inputs often necessitates multi-view data or costly per-object optimization to resolve ambiguities. The second category targets direct, diffusion-based 2D relighting [11, 52, 56, 28]. Methods like DiLightNet and IC-Light exploit diffusion priors to produce high-fidelity relit images with fine-grained control. However, these approaches are typically computationally expensive, stochastic, and operate in 2D, failing to guarantee multi-view consistency for 3D applications.
@@ 2.2 Generative 3D Priors and Representations
Diffusion priors and score-distillation sampling (SDS) have catalyzed rapid progress in text-to-3D and image-to-3D generation [34, 40, 47, 38]. While SDS-based methods transfer 2D generative knowledge to 3D effectively, they suffer from slow iterative optimization. Consequently, recent works have shifted toward feed-forward 3D reconstruction models trained on large-scale 3D datasets [14, 45, 55]. Specifically, Trellis [45] utilizes Structured 3D Latents (SLAT) to compress complex geometry and appearance into sparse tokens, enabling efficient decoding.
Concurrently, 3D Gaussian Splatting (3DGS) [18] has emerged as a rasterization-friendly representation supporting real-time differentiable rendering. While current feed-forward models (like LRM or Trellis) excel at geometry, they typically bake lighting into the texture, limiting downstream utility. Our method builds upon the efficiency of SLAT and 3DGS but fundamentally redesigns the generation process. We introduce a lighting-homogenized variant of SLAT and a custom neural decoder, replacing static texture prediction with a relightable neural field.
@@ 2.3 Relightable 3D asset synthesis
Producing relightable 3D assets requires models to represent both intrinsic surface properties and lighting-dependent transport (shadows, speculars, interreflections). Prior works condition NeRFs, Gaussian splats or meshes on lighting inputs to enable relighting-aware outputs [51, 16, 36, 21, 12, 46, 2, 59, 33]. Many approaches either use volumetric neural renderers that are costly at inference, or attempt to estimate PBR maps without lighting supervision, which leads to poor disentanglement [35, 26, 39]. Some models explore large inverse-rendering architectures to predict PBR properties from sparse views, but computational cost and optimization per-object remain bottlenecks [22, 58]. Recent works [10, 41] employ diffusion models to generate multi-view material maps or multi-view relighted images, followed by 3D reconstruction. However, the absence of explicit 3D constraints in the generation stage makes it difficult to guarantee consistency across views.
@@
Figure 3: Pipeline of NeAR. Stage 1: Lighting Homogenization extracts a Lighting-Homogenized Structured 3D Latent (LH-SLAT) from a casually lit input image. Stage 2: Relightable 3DGS Synthesis generates a relightable 3D Gaussian Splatting (3DGS) field conditioned on the LH-SLAT, target illumination, and target viewpoint. The decoded 3DGS encodes geometry, appearance, and light–material interactions, and is rendered into the final relit image.
@@ 3.1 Preliminary
3D Gaussian Splatting (3DGS). 3DGS [18] represents scenes with anisotropic Gaussians, rendered via splatting and $\alpha$-blending: $C=\sum_{i\in\mathcal{N}}c_{i}\alpha_{i}\prod_{j<i}(1-\alpha_{j})$. Crucially, standard 3DGS models color $c_{i}$ using Spherical Harmonics (SH), which inherently bakes static lighting into the representation. To enable relighting, we forego SH and predict color dynamically conditioned on target illumination.
Structured 3D Latents (SLAT). Following Trellis [45], we use SLAT to encode 3D assets efficiently. A SLAT $\mathcal{Z}=\{(\bm{z}_{k},\bm{p}_{k})\}_{k=1}^{K}$ consists of $K$ active feature tokens, where each token $\bm{z}_{k}\in\mathbb{R}^{D}$ is associated with a coordinate $\bm{p}_{k}$ in a sparse voxel grid. This representation focuses capacity on surface regions ($K\ll N^{3}$) and supports diverse decoding heads. However, standard SLATs blindly encode input appearance—including shadows and highlights. Our goal is to transform $\mathcal{Z}$ into a lighting-homogenized form, canonicalizing the appearance to a uniform illumination while preserving geometry.
@@ 3.2 Overview of NeAR
The challenge in single-image 3D relighting lies in disentangling lighting from intrinsic object properties, since shadows, highlights, and interreflections are inherently entangled with geometry. To avoid unstable PBR inversion and black-box neural generation, we propose a homogenize-then-synthesize framework. NeAR first extracts a Lighting-Homogenized SLAT (LH-SLAT) to neutralize lighting effects, then decodes a relightable 3DGS. Our framework consists of two stages:
Stage 1: Lighting Homogenization. We first utilize the pre-trained flow model $f_{s}$ to map the arbitrarily lit input $I_{\text{in}}$ into an initial shaded SLAT $Z_{s}$. Operating within this sparse voxel space, we employ a LoRA-adapted model $f_{\theta}$ to steer the latent representation from $Z_{s}$ toward a Lighting-Homogenized SLAT (LH-SLAT) $Z_{\text{lh}}$:
@@
Stage 2: Relightable 3DGS Synthesis. Leveraging the homogenized representation $Z_{\text{lh}}$, a feed-forward decoder $\mathcal{D}$ synthesizes a relightable Gaussian field $\mathcal{G}$. This process is conditioned on the target view $\mathbf{v}_{\text{target}}$ and the target illumination $L_{\text{target}}$, encoded via $\mathcal{E}_{l}$:
@@
3.3 LH-SLAT Reconstruction & Generation
@@
LH-SLAT Generation. As shown in 3 top right corner, we use a rectified flow model $f_{\theta}$ to generate the lighting-homogenized SLAT $Z_{\text{lh}}$ from the input image $I_{\text{in}}$. The rectified flow model is trained to learn the mapping between the arbitrarily lit image and the corresponding latent representation under our homogenized lighting conditions. Specifically, we utilize a pre-trained SLAT rectified flow model $f_{s}$ to generate the shadowed SLAT $Z_{s}$ from the input image $I_{\text{in}}$, and subsequently fine-tune $f_{s}$ using LoRA  [15] in the sparse voxel space  [45] to achieve lighting homogenization. The loss function for training is the conditional flow matching loss $\mathcal{L}_{stage1}$:

### v2

@@ Abstract
Neural asset authoring and neural rendering have traditionally evolved as disjoint paradigms: one generates digital assets for fixed graphics pipelines, while the other maps conventional assets to images. However, treating them as independent entities limits the potential for end-to-end optimization in fidelity and consistency. In this paper, we bridge this gap with NeAR, a Coupled Neural Asset–Renderer Stack. We argue that co-designing the asset representation and the renderer creates a robust ”contract” for superior generation. On the asset side, we introduce the Lighting-Homogenized SLAT (LH-SLAT). Leveraging a rectified-flow model, NeAR lifts casually lit single images into a canonical, illumination-invariant latent space, effectively suppressing baked-in shadows and highlights. On the renderer side, we design a lighting-aware neural decoder tailored to interpret these homogenized latents. Conditioned on HDR environment maps and camera views, it synthesizes relightable 3D Gaussian splats in real-time without per-object optimization. We validate NeAR on four tasks: (1) G-buffer-based forward rendering, (2) random-lit reconstruction, (3) unknown-lit relighting, and (4) novel-view relighting. Extensive experiments demonstrate that our coupled stack outperforms state-of-the-art baselines in both quantitative metrics and perceptual quality. We hope this coupled asset-renderer perspective inspires future graphics stacks that view neural assets and renderers as co-designed components instead of independent entities.
Figure 1: Comparison of NeAR and Decoupled Paradigms. Left: Visual results under target illumination. Cols. 3–5 are rendered via Blender to evaluate asset quality. Insets (right of cols. 4&5) display PBR maps (top-down: Base Color, Metallic, Roughness). Baselines suffer from baked-in lighting (Trellis) or material ambiguity (HY3D-2.1). Notably, HY3D-2.1 wrongly assigns high metallic values to the bread (see Metallic map, Row 1) and exhibits inconsistent highlights on the robot (Row 3). While our intermediate PBR decomposition (col. 5) corrects materials, it struggles with complex effects like transparency (Helmet, Row 2) under standard rendering. Our full Neural Renderer (col. 6) resolves this, yielding photorealistic results closest to GT. Right: Quantitative results on the Glossy Synthetic dataset. NeAR achieves the highest PSNR across all four tasks, demonstrating the superiority of our coupled stack.
@@ 1 Introduction
Images are determined by the interaction of light with scene geometry, materials, and lighting. Classical computer graphics separates this process into asset authoring, where artists define scene properties, and rendering, where a physically based renderer simulates light transport. While effective, this separation requires substantial manual effort, computationally expensive simulations, and makes inverse reconstruction from real-world images or video challenging. Recent advances in neural graphics [zhang2024clay, wu2024direct3d, li2025triposg, zhao2025hunyuan3d, xiang2024structured, ye2025hi3dgen, sf3d, chen2025meshgen, jin2024neural, rgbx, zeng2024dilightnet, magar2025lightlab] address these limitations from two complementary directions: neural asset authoring uses generative models [zhang2024clay, wu2024direct3d, li2025triposg, zhao2025hunyuan3d, xiang2024structured, ye2025hi3dgen, chen20243dtopia, sf3d, chen2025meshgen] to synthesize full 3D assets for traditional pipelines, reducing manual effort, while neural renderers map these assets—often converted into intermediate representations such as depth, normals, or shading buffers—directly to images [diffusionrenderer, rgbx, zeng2024dilightnet], providing a data-driven alternative to analytic rendering and enabling more robust inverse inference.
@@
Despite recent progress in generating 3D assets with PBR materials [zhang2024clay, wu2024direct3d, li2025triposg, zhao2025hunyuan3d, xiang2024structured, ye2025hi3dgen, chen20243dtopia, sf3d, chen2025meshgen], a fundamental limitation remains: asset generation and neural rendering are typically developed in isolation, with assets created assuming a fixed renderer and renderers trained on static asset distributions. This separation becomes problematic when errors in asset decomposition—such as misidentified albedo or incorrect normal maps—propagate through the rendering pipeline. Because rendering is a nonlinear process, small errors in asset decomposition compound into visible artifacts like baked-in shadows or lighting inconsistencies. Fig. 1 demonstrates this issue: existing methods rendered with traditional physically-based renderers (e.g., Blender) exhibit lighting artifacts and fail to achieve faithful relighting.
@@
To this end, we propose NeAR, a Coupled Neural Asset–Renderer Stack for single-image relightable 3D generation. Our key insight is to co-design the asset representation and rendering process to enable relighting directly through a shared, lighting-homogenized latent space. On the asset side, we introduce a Lighting-Homogenized Structured 3D Latent (LH-SLAT). Unlike standard assets that rely on fragile explicit decomposition, our model lifts the casually lit input into a canonical latent form. As visualized in Fig. 2, this process transforms a shadow-affected representation (Shaded-SLAT) into a clean, homogenized state, effectively suppressing baked-in shadows and unstable highlights while preserving geometric cues. On the renderer side, we design a lighting-aware neural renderer. Conditioned on a lighting tokenizer, this renderer learns to interpret the homogenized latents and synthesize view-dependent appearance under arbitrary HDR environments via differentiable 3D Gaussian splatting. By unifying the representation, NeAR generates assets that naturally support real-time, high-quality relighting and novel-view synthesis with consistent materials across views. Fig. 3 shows a comparison between our method and previous single-image relighting frameworks.
Figure 2: Lighting homogenization as the bridge between assets and renderer. We visualize the intrinsic components (Base Color, Ambient Occlusion), rendering results under random and uniform lighting, shadow maps, as well as relighting outputs generated respectively by Shaded SLAT and LH-SLAT. By mapping casually lit images to a canonical illumination space, LH-SLAT effectively suppresses baked-in shadows and unstable specularities while preserving geometry-consistent diffuse cues. This stable latent space serves as the robust ”contract” for our lighting-aware neural renderer to enable controllable relighting.
Figure 3: Overview of NeAR vs. Existing Frameworks. (a-b) Existing 2D methods lack explicit 3D awareness; specifically, (a) struggles to disentangle specular highlights, while both fail to guarantee multi-view consistency during relighting. (c) State-of-the-art 3D generation methods decouple asset authoring from rendering, relying on ill-posed PBR decomposition that often results in material inaccuracies and baked-in artifacts. In contrast, (d) NeAR (Ours) employs a Coupled Neural Asset–Renderer Stack. By utilizing the LH-SLAT representation, we simultaneously achieve photorealistic relighting and consistent novel-view synthesis.
@@
    Coupled neural asset–renderer stack. We introduce NeAR, an learnable graphics stack where the neural asset representation and neural renderer are co-designed for single-image relightable 3D asset generation.
@@ 2.1 Image relighting and inverse rendering
Image relighting and inverse rendering lie at the intersection of geometry, material estimation, and light transport and have been studied from both physics-driven and data-driven perspectives [jin2024neural]. Classical inverse rendering methods (e.g., SIRFS) recover interpretable PBR maps (albedo, roughness, normals) via optimization and hand-crafted priors [barron2013intrinsic]. These pipelines are interpretable and editable, but the inverse problem is highly ill-posed in real scenes: shadows, inter-reflections and view-dependent highlights easily bias the recovered materials and produce baked-in artifacts under re-rendering.
Recent learning-based approaches generally fall into two categories. The first focuses on physically-structured decomposition [zhao2025hunyuan3d, diffusionrenderer, engelhardt2025svim3d]. Although decomposition yields interpretable assets, accurate regression from casual single-view inputs often necessitates multi-view data or costly per-object optimization to resolve ambiguities. The second category targets direct, diffusion-based 2D relighting [fortier2024spotlight, zeng2024dilightnet, zhang2025scaling, magar2025lightlab]. Methods like DiLightNet and IC-Light exploit diffusion priors to produce high-fidelity relit images with fine-grained control. However, these approaches are typically computationally expensive, stochastic, and operate in 2D, failing to guarantee multi-view consistency for 3D applications.
@@ 2.2 Generative 3D Priors and Representations
Diffusion priors and score-distillation sampling (SDS) have catalyzed rapid progress in text-to-3D and image-to-3D generation [poole2022dreamfusion, tang2023dreamgaussian, yan2025learning, shi2023mvdream]. While SDS-based methods transfer 2D generative knowledge to 3D effectively, they suffer from slow iterative optimization. Consequently, recent works have shifted toward feed-forward 3D reconstruction models trained on large-scale 3D datasets [hong2023lrm, xiang2024structured, zhang2024clay]. Specifically, Trellis [xiang2024structured] utilizes Structured 3D Latents (SLAT) to compress complex geometry and appearance into sparse tokens, enabling efficient decoding.
Concurrently, 3D Gaussian Splatting (3DGS) [kerbl20233d] has emerged as a rasterization-friendly representation supporting real-time differentiable rendering. While current feed-forward models (like LRM or Trellis) excel at geometry, they typically bake lighting into the texture, limiting downstream utility. Our method builds upon the efficiency of SLAT and 3DGS but fundamentally redesigns the generation process. We introduce a lighting-homogenized variant of SLAT and a custom neural decoder, replacing static texture prediction with a relightable neural field.
@@ 2.3 Relightable 3D asset synthesis
Producing relightable 3D assets requires models to represent both intrinsic surface properties and lighting-dependent transport (shadows, speculars, interreflections). Prior works condition NeRFs, Gaussian splats or meshes on lighting inputs to enable relighting-aware outputs [zeng2023relighting, jin2024neural, remondino2023critical, li2023relit, gao2024relightable, yan20243dsceneeditor, bi2024gs3, zhao2024illuminerf, Poirier_Ginter_2024]. Many approaches either use volumetric neural renderers that are costly at inference, or attempt to estimate PBR maps without lighting supervision, which leads to poor disentanglement [qiu2024richdreamer, liu2024unidream, shim2024mvlight]. Some models explore large inverse-rendering architectures to predict PBR properties from sparse views, but computational cost and optimization per-object remain bottlenecks [li2025lirm, zhang2024relitlrm]. Recent works [engelhardt2025svim3d, tang2025rogr] employ diffusion models to generate multi-view material maps or multi-view relighted images, followed by 3D reconstruction. However, the absence of explicit 3D constraints in the generation stage makes it difficult to guarantee consistency across views.
@@
Figure 4: Pipeline of NeAR as a coupled neural asset-renderer stack. Top (Inference Stage): An end-to-end inference pipeline. Given a single image and a geometry prior (e.g., mesh from HY3D), Stage 1 utilizes a rectified-flow backbone with LoRA adaptation to predict the Lighting-Homogenized SLAT (LH-SLAT). This latent acts as a bridge, which is then consumed by the Stage 2 lighting-aware neural renderer to synthesize relightable 3DGS under novel illumination and viewpoints. Bottom-Left (Data Prep): Offline construction of ground-truth LH-SLATs by rendering assets under homogenized illumination and encoding them via a sparse VAE. Bottom-Right (GS Decoding & Rendering): Detailed architecture of the 3DGS decoding head, which predicts Gaussian attributes from lighting-dependent features, followed by a differentiable rasterizer $\mathcal{M}$ that renders the final HDR image, shadow and PBR auxiliary maps.
@@ 3.1 Preliminary
3D Gaussian Splatting (3DGS). 3DGS [kerbl20233d] represents scenes with anisotropic Gaussians, rendered via splatting and $\alpha$-blending: $C=\sum_{i\in\mathcal{N}}c_{i}\alpha_{i}\prod_{j<i}(1-\alpha_{j})$. Crucially, standard 3DGS models color $c_{i}$ using Spherical Harmonics (SH), which inherently bakes static lighting into the representation. To enable relighting, we forego SH and predict color dynamically conditioned on target illumination.
Structured 3D Latents (SLAT). Following Trellis [xiang2024structured], we use SLAT to encode 3D assets efficiently. A SLAT $\mathcal{Z}=\{(\bm{z}_{k},\bm{p}_{k})\}_{k=1}^{K}$ consists of $K$ active feature tokens, where each token $\bm{z}_{k}\in\mathbb{R}^{D}$ is associated with a coordinate $\bm{p}_{k}$ in a sparse voxel grid. This representation focuses capacity on surface regions ($K\ll N^{3}$) and supports diverse decoding heads. However, standard SLATs blindly encode input appearance—including shadows and highlights. Our goal is to transform $\mathcal{Z}$ into a lighting-homogenized form, canonicalizing the appearance to a uniform illumination while preserving geometry.
@@ 3.2 Overview of NeAR
The challenge in single-image 3D relighting lies in disentangling lighting from intrinsic object properties, since shadows, highlights, and interreflections are inherently entangled with geometry. To avoid unstable PBR inversion and black-box neural generation, we propose a homogenize-then-synthesize framework that functions as a coupled stack. NeAR first extracts a Lighting-Homogenized SLAT (LH-SLAT) from the input image to neutralize lighting effects, then decodes a relightable 3DGS. Our framework consists of two stages:
Stage 1: Light Homogenization-SLAT Generation. We first utilize the pre-trained flow model $f_{s}$ to map the arbitrarily lit input $I_{\text{in}}$ into an initial shaded SLAT $Z_{s}$. Operating within this sparse voxel space, we employ a LoRA-adapted model $f_{\theta}$ to steer the latent representation from $Z_{s}$ toward a Lighting-Homogenized SLAT (LH-SLAT) $Z_{\text{lh}}$:
@@
Stage 2: Relightable Neural 3DGS Synthesis. Leveraging the homogenized representation $Z_{\text{lh}}$, a feed-forward decoder $\mathcal{D}$ synthesizes a relightable Gaussian field $\mathcal{G}$. This process is conditioned on the target view $\mathbf{v}_{\text{target}}$ and the target illumination $L_{\text{target}}$, encoded via $\mathcal{E}_{l}$:
@@
3.3 Light Homogenization & LH-SLAT Rec.
@@
3.4 LH-SLAT Generation
As shown in Fig. 4 top right corner, we use a rectified flow model $f_{\theta}$ to generate the lighting-homogenized SLAT $Z_{\text{lh}}$ from the input image $I_{\text{in}}$. The rectified flow model is trained to learn the mapping between the arbitrarily lit image and the corresponding latent representation under our homogenized lighting conditions. Specifically, we utilize a pre-trained SLAT rectified flow model $f_{s}$ to generate the shadowed SLAT $Z_{s}$ from the input image $I_{\text{in}}$, and subsequently fine-tune $f_{s}$ using LoRA  [hu2022lora] in the sparse voxel space  [xiang2024structured] to achieve lighting homogenization. The loss function for training is the conditional flow matching loss $\mathcal{L}_{stage1}$: